Grade 11Algebra

if z=x+iy and |2z+1|=|Z-2i| show that 3(x2+y2)+4(x+y)=3
X2 IS X SQUARE
Y2 IS Y SQUARE

subhajit sardar
6 Years agoGrade 11

1 Answer

Arun
6 Years ago
Dear student
 
 
Dear student 
 
|2x + 1 +2 iy| = |x + i(y – 2)|
 
4x^2 + 1 + 4x +4 y^2 = x^2 + y^2 + 4 - 4y
 
3 x^2 + 3 y^2 + 4x + 4y  = 3
 
Hope it helps